A gaseous hydrocarbon contains 85.7% carbon and 14.3% hydrogen. 1 litre of the hydrocarbon weighs 1.26 g at NTP. Determine the Molecular Formula of the

Grade:11

1 Answers

Vikas TU
7 years ago
Let the hydrocarbon be CxHy its molar mass = 12x + y
moles of C = 85.7/12x
moles of H = 14.3/y
 
At NTP, moles of hydrocarbon = 1/22.4
molar mass = moles * 1.26
 = 1.26/22.4
= 0.05625
 
12x + y = 0.05625......................................................(A)
1.26/(12x + y)  = 85.7/12x + 14.3/y...............................(B)
Solve them get x and y.